Database last updated on Mon, 02 Feb 2026 01:10:59 +0000 by localhost
Critical warnings from b-files' comments for days between 2024001 and 2024001 (0 sec, 3 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| hp: | Spectrometer alignment trouble. | 5 | 1 | B00124.187 | TORON | #187 | 13:21:38 | | 1 | -1 |
| sl: | Lamp not on... sl test terminated F1= 0 F4= 0 | 1 | 1 | B00124.187 | TORON | #187 | 14:21:07 | | 1 | -1 |
| re: | Previous RE didn't seem to help, so not doing reset again. Operator intervention required. Quitting to menu. | 1 | 1 | B00124.187 | TORON | #187 | 14:39:05 | | 1 | -1 |
Instrument constant changes for the week before 2024001 (2 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 10 rows)
HP with step corrections larger than 20 steps (0 sec, 1 rows)
| Serial | YYYYJJJ | HP large | max correction |
| 187 | 2024001 | 1 | 79.9 |
Non-critical warnings from b-files' comments for days between 2024001 and 2024001 (0 sec, 1 rows)
| Type | Comment | count | priority | fname | location | serial | time | value | min | max |
| pd: | Going to menu | 1 | 99 | B00124.187 | TORON | #187 | 14:39:06 | | 1 | -1 |
Measurements and tests summary.
| _jday | _loc | _serial | _skc | dt | fm | hg | hp | rs | sl | uv |
| 001 | TORONTO | #187 | ecsc | 1 | 35 | 6 | 8 | 1 | 2 | 11 |
(0 sec)
Relative and absolute humidity for day 2024001 (1 sec, 1 rows)
| Serial | RH, % | AH, g/m3 | 7-day RH history | Des-nt replcd days ago |
| 187 | 4.2 | 0.4 |